OpenStudy (anonymous):

Rocket-powered sleds are been used to test the responses of humans to acceleration. Start- ing from rest, one sled can reach a speed of 530 m/s in 1.58 s and can be brought to a stop again in 2.7 s. Find the acceleration of the sled while ac- celerating. Answer in units of m/s2 Find the acceleration of the sled when brak- ing. Answer in units of m/s2

OpenStudy (agent0smith):

acceleration = change in speed / time

OpenStudy (agent0smith):

speed of 530 m/s in 1.58 s = change in speed is 530m/s can be brought to a stop again in 2.7 s. = change in speed is -530m/s (from 530 m/s to zero m/s)

OpenStudy (john_es):

Yes it is that way. \[a=\frac{v_f-v_0}{\Delta t}\] In the first case, v0=0 m/s, vf=530 m/s, Δt=1.58 s. In the second, vf=0 m/s, v0=530 m/s, Δt=2.7 s

OpenStudy (anonymous):

hmm so it's just 335.443 m/s^2 and 196.296m/s^2

OpenStudy (anonymous):

the 196.296 answer is wrong... do you know why?

OpenStudy (anonymous):

Oh wait is it negative?

OpenStudy (john_es):

Yes! Because you doesn't put the - sign.

OpenStudy (john_es):

The value must be negative.

OpenStudy (john_es):

;)

OpenStudy (anonymous):

whoops forgot that was important ._. thanks

OpenStudy (john_es):

You're welcome.

OpenStudy (agent0smith):

335.443 m/s^2 damn, that's like 33+ g's, for 1.5 seconds